# Systematically Improvable Numerical Atomic Orbital Basis Using Contracted Truncated Spherical Waves

### 14 Mar 2026 v1

@@ Abstract
To solve the Kohn-Sham equation within the framework of density functional theory, we develop a scheme to construct numerical atomic orbital (NAO) basis sets by contracting truncated spherical waves (TSWs). The contraction minimizes the trace of the kinetic operator in the residual space, generalizing the spillage minimizing scheme [M. Chen et al., J. Phys. Condens. Matter 22, 445501 (2010); P. Lin et al., Phys. Rev. B 103, 235131 (2021)]. In addition to the systematic improvability inherited from previous schemes, the use of TSW instead of plane waves as the expansion basis bridges reference states and NAOs more effectively, and eliminates spurious interactions between periodic images, thereby enabling better transferability through the inclusion of extensive reference states. Benchmarks demonstrate that the constructed NAO achieves satisfactory precision for various properties of both molecules and bulk systems, including total energy, bond length, atomization energy, lattice constant, cohesive energy, band gap, and energy-level alignment. By incorporating unoccupied states, the improved transferability in describing the conduction band is demonstrated to be effective and substantial.
@@ I Introduction
Atomic orbital basis strikes a balance between computational efficiency and accuracy in the description of the electronic structure of materials. Nowadays, well-developed atomic orbital basis sets can achieve meV-level energy convergence across a variety of chemical environments [11]. In the meantime, their numbers of basis functions are usually quite small compared to other options, and computational cost might be further reduced by exploiting their locality to enable linear-scaling methods [20, 3]. Because of these advantages, they are widely used in systems of various sizes, ranging from a few to thousands of atoms [57, 37, 70, 48, 2, 23, 47, 43, 22, 87, 106, 6, 27, 1, 99].
@@
Despite those advantages, the other side of the coin is that the atomic orbital basis lacks the systematic convergence possessed by many other basis types. For example, the completeness of the plane-wave basis can be entirely controlled by the kinetic energy cutoff. Even within the family of localized basis, there are also systematically refinable options like real space grids [7], truncated spherical waves [31], B-splines [32], Lagrange functions [101], wavelets [24], etc. While this fact poses no fundamental difficulty, additional effort is required to generate a sensible hierarchy of basis sets with transferable accuracy.
@@
In the past three decades, much interest has been devoted to the development of NAOs, whose radial functions are numerically tabulated and are fully flexible within a cutoff radius. The strategy for generating strictly localized NAOs is not unique. For example, a wide range of methods involve finding the Kohn-Sham orbitals of isolated atoms subject to some confining potential (so-called pseudo-atomic orbitals, PAOs) [85, 79, 35, 41, 89, 4]. Based on this idea, Ozaki [73, 72] proposes that PAOs be used as “primitive functions” to construct NAOs, where the coefficients are optimized alongside self-consistent cycles (of course, the resulting NAOs are closely related to the chemical environment and are not basis sets in the conventional sense). Alternatively, Blum et al [11] use the occupied PAOs as the minimal basis set, and enlarge it by iteratively picking up the function out of a pool of predefined candidates - including cation-like and hydrogen-like radial functions - that best improve some target energy, like the non-self-consistent total energies of several selected dimers. Furthermore, the concept of “correlation consistency” proposed by Dunning [19] is also explored within the context of NAO by Zhang et al [111]. By minimizing the frozen-core RPA@PBE total energies of free atoms, the resulting basis sets are suitable for correlation methods that include explicit sums over unoccupied states, such as RPA or MP2.
@@
Methods that adopt strategies other than explicit energy optimization also exist. Sánchez-Portal et al [84, 83] proposed to generate NAOs towards some reference states by optimizing the “spillage”. In their original scheme, the primitive functions that constitute NAOs are PAOs or Slater-type orbitals, and the reference systems are ordinary solids. Following their work, Kenny et al [44] used confined neutral/charged atoms instead of solids as reference systems. Later, Chen, Guo, and He (CGH)  [15] proposed using truncated spherical waves (TSW) as primitive functions to gain greater flexibility and a series of dimers of variable bond lengths as reference systems to improve transferability. Recently, Lin, Ren, and He (LRH)  [53] show that CGH basis sets can be further improved by introducing an extra gradient term to the original spillage function. All NAOs developed above have a striking balance of efficiency and precision, enabling large-scale and accurate simulations in various fields of research [47, 23, 78, 90, 68, 14, 62, 77, 57, 55, 13, 49, 69, 46, 67, 71, 74, 91, 95, 38, 50, 9, 21, 52, 82, 59, 58, 60, 92, 39, 112, 113, 64, 94].
Although LRH basis sets can produce remarkable accuracy in the calculations of structural and electronic properties for a number of molecules and bulk materials, there is still room for improvement on this type of basis set. The hierarchy of LRH basis sets is built upon split-valence and polarization, where polarization orbitals are restricted to have an angular momentum merely one larger than that of the valence orbitals. The lack of involving orbitals with higher angular momentum clearly hinders the completeness: in Figure. 1 of their original work, The energy difference between the plane wave (PW) and TSWs (with restricted angular momentum) can reach tens of meV. Given that TSW are systematically improvable towards a complete basis, such a gap might be significantly improved [8]. In addition, while including more reference states improves the transferability of NAOs, we find the spurious interactions with the periodic images that causes the artificial bonding across the vacuum is hard to avoid with PW as expansion basis in practice (Appendix A), though PW is usually a feasible choice to approach the complete basis set (CBS).
In this work, we propose an NAOs generation scheme by directly contracting the near-complete TSW. Our NAOs not only inherit the merits of those built with previous schemes for precise energy calculations and structural property predictions, but also achieve improved accuracy in tasks involving unoccupied states, such as absorption spectra via time-dependent DFT (TD-DFT).
@@ II.1 Parametrization
The parametrization of the radial shape typically balances several factors: computational efficiency of integrals, locality, and flexibility. An ideal form should have strict locality and maximum flexibility within some reasonable cutoff radius while enabling efficient evaluation of common integrals.
Over the past few decades, many numerical algorithms have been proposed [93, 88, 97, 30] to perform integrations for NAOs efficiently, and the computational efficiency has become less of a concern. For strict locality and maximum flexibility, TSWs [31] stand out as a sensible choice as the building blocks of the atomic orbitals, which gives
$\chi_{l\zeta}(r)=\left\{\begin{matrix}\displaystyle\sum_{q=1}^{N_{l}}j_{l}(\theta_{lq}r/r_{c})c_{lq\zeta}&r\leq r_{c}\\[6.0pt] 0&r>r_{c}\end{matrix}\right.$~~. (1)
Here $j_{l}$ is the spherical Bessel function of the first kind, $\theta_{lq}$ is the $q$-th positive zero of $j_{l}$, and $r_{c}$ is the cutoff radius. Note that spherical waves are also eigenfunctions of the kinetic energy operator (assuming atomic Rydberg unit)
@@
so, similar to the plane wave basis, the number of spherical waves ($N_{l}$) can be controlled by a kinetic energy cutoff
$N_{l}=\text{max}\{q|\theta_{lq}<r_{c}\sqrt{E_{c}}\}~~.$ (3)
The set of all TSWs approaches CBS as the cutoff radius, maximum angular momentum, and cutoff energy all approach infinity. Given a cutoff radius, the task of building an atomic orbital basis can be viewed as seeking the contraction coefficients of TSWs.
Note that TSWs themselves do not have vanishing derivatives at the cutoff radius. To avoid discontinuities of derivatives, several previous works [44, 15] applied an extra “smoothing function”. Though having no significant effect on the quality or implementation as NAO, such a modification not only demands some extra effort into finding out an optimal smoothing parameter on top of the optimization of contraction coefficients, but also makes orbitals no longer pure combinations of TSWs, thereby losing the possibility of benefiting from certain analytical expressions of integrals [31, 66].
Instead of modifying the form, we suggest seeking the subspace of zero derivatives at the cutoff while preserving the form. Suppose we would like to combine spherical Bessel functions into a new set of basis functions with vanishing first $M$ derivatives at the cutoff
$\xi_{l\lambda}(r)$ $=\sum_{q=1}^{N_{l}}j_{l}(\theta_{lq}r/r_{c})K_{q\lambda}~~,$ (4)
@@
Eqs. 4-6 gives $DK=0$, i.e., the columns of $K$ reside in the null space of $D$. In practice, they can be taken as the set of right singular vectors corresponding to zero singular values. Moreover, since $j_{l}$ is the solution to the spherical Bessel equation
@@
we have $D_{2q}=(-2/r_{c})D_{1q}$. In other words, any linear combination that eliminates the first derivative also eliminates the second derivative. If only the first two derivatives are concerned, we can have a set of $N_{l}-1$ basis functions $\{\xi_{l\lambda}\}$ with vanishing first two derivatives out of $N_{l}$ spherical Bessel functions. In the rest of this paper, we denote $\{\xi_{l\lambda}\}$ as NSW and unless otherwise noted, orbitals are made by NSW with $M=2$.
@@ II.2 Optimization
In this section, the contraction of the NSW basis set to obtain optimized NAOs is demonstrated in terms of defining the generalized spillage, choosing reference systems, and the NAO basis set hierarchy. The summarized workflow is shown in Fig. 1.
@@ II.2.1 Generalized Spillage
proposed by Lin et al[53] ($\hat{p}$ is the momentum operator), we define the function $\tilde{\mathcal{S}}$ to minimize as
@@
This expression describes a different picture from Sánchez-Portal et al., Chen et al. (eqn. 8) and Lin et al. (eqn.10): eqn. 11 corresponds to the trace of any arbitrary operator $\hat{O}$ in the residual space spanned by $\{(1-\hat{P}_{\mathbf{k}})|\psi_{n\mathbf{k}}^{\mathrm{NSW}}\rangle\}$. $\{|\psi_{n\mathbf{k}}^{\mathrm{NSW}}\rangle\}$ is some set of vectors in space spanned by NSWs, and the subspace spanned by contracted NSWs (NAOs) is found by minimizing the residual information defined by a given operator as much as possible. In this paper, $\{|\psi_{n\mathbf{k}}^{\mathrm{NSW}}\rangle\}$ is taken to be the occupied bands optionally plus a subset of unoccupied bands, and $\hat{O}$ is taken to be the kinetic operator $\hat{p}^{2}$. Comparatively, if $|\psi_{n\mathbf{k}}\rangle$ is taken to be the reference states with PW basis, substituting $\hat{O}$ with the overlap $\hat{S}$ or overlap plus kinetic operator $\hat{S}+\hat{T}$ will reduce eqn. 11 to eqn. 8 or eqn. 10, respectively. Given that the space spanned by TSWs or NSWs is not strictly the subspace of PWs because of the real-space truncation on Bessel functions at $r_{\mathrm{c}}$, minimizing spillage defines a fitting problem instead.
@@ II.2.2 Reference Systems
For the transferability of NAOs, reference systems should cover potential polarization or bonding in real chemical environments. Blum et al[11] and Chen et al[15] use homo-nuclear dimers with various bond lengths, and Lin et al[53] include additional trimers in triple-zeta level NAOs generation. All their NAOs show stable precision across a wide range of systems. In this paper, we include at least 4 bond lengths for each dimer as reference. The detailed values of bond lengths are selected such that the corresponding energies relative to the equilibrium are within an energy window $\sim$1.0 eV.
@@ II.3 Basis Set Hierarchy
To hierarchically approach the maximal angular momentum of NSW, we employ Dunning’s hierarchy [19] to construct numerical atomic orbitals (NAOs). We construct three standard NAO tiers: minimal, polarized valence double-zeta (pVDZ), and polarized valence triple-zeta (pVTZ), along with two variants with restricted angular momentum, denoted pVTZ- and pVQZ=, for comparison.
Our approach adopts the ”shell-wise” optimization scheme established in the CGH [15] and LRH [53] basis sets, wherein basis functions are incrementally added and optimized while keeping all previously generated orbitals fixed. The procedure begins by generating a minimal basis corresponding to the pseudopotential valence-electron configuration. New functions are then added and optimized with the minimal basis held fixed to form the pVDZ set. Subsequent steps involve adding further functions under fixed pVDZ constraints to produce the angular-momentum-restricted triple-zeta basis pVTZ-. Finally, additional extensions yield the full triple-zeta pVTZ and the doubly restricted quadruple-zeta pVQZ= basis sets.
@@
Figure 1: The workflow of NAOs generation. The total workflow is divided into 4 steps. In the first step, a PW basis set is prepared, and a set of structures is chosen. In the second step, the superparameters $r_{\mathrm{c}}$ and $l_{\max}$ are chosen so that the energy difference between PW and NSW $\epsilon^{\mathrm{PW}}_{\mathrm{NSW}}$ can be smaller than a given threshold. In the third step, series of SCF calculations are performed with converged NSW, yield the overlap $S_{\mu\nu}(\mathbf{k})$, kinetic energy $T_{\mu\nu}(\mathbf{k})$ matrices, and wavefunction $C_{n\mu}(\mathbf{k})$, in which $\mu$ and $\nu$ run over all NSW, $n$ denotes eigenstates, and $\mathbf{k}$ distinguishes different $\mathbf{k}$-points. In the last step, the generalized spillage (eqn. 11) is minimized, during which the contraction of NSW is done.

### 09 Apr 2026 v2

@@ Abstract
To solve the Kohn-Sham equation within the framework of density functional theory, we develop a scheme to construct numerical atomic orbital (NAO) basis sets by contracting truncated spherical waves (TSWs). The contraction minimizes the trace of the kinetic operator in the residual space, generalizing the spillage minimizing scheme [M. Chen et al., J. Phys. Condens. Matter 22, 445501 (2010); P. Lin et al., Phys. Rev. B 103, 235131 (2021)]. In addition to the systematic improvability inherited from previous schemes, the use of TSWs instead of plane waves as the expansion basis bridges reference states and NAOs more effectively, and eliminates spurious interactions between periodic images, thereby enabling better transferability through the inclusion of extensive reference states. Benchmarks demonstrate that the constructed NAO achieves satisfactory precision for various properties of both molecules and bulk systems, including total energy, bond length, atomization energy, lattice constant, cohesive energy, band gap, and energy-level alignment. Furthermore, by incorporating unoccupied states, the improved transferability in describing the conduction band is demonstrated to be effective and substantial.
@@ I Introduction
Density functional theory (DFT) [34, 48] has found extensive application in computational investigations in materials science. Atomic orbital basis strikes a balance between computational efficiency and accuracy in the description of the electronic structure of materials. Nowadays, well-developed atomic orbital basis sets can achieve meV-level energy convergence across a variety of chemical environments [11]. In general, the number of atomic orbital basis sets is smaller than that of alternative options, and computational costs can be further reduced by leveraging their locality to implement linear-scaling methods. [20, 3]. Owing to these inherent advantages, they have been widely employed for atomic systems of various sizes, ranging from a few atoms up to thousands of atoms. [59, 38, 72, 50, 2, 23, 49, 44, 22, 89, 108, 6, 27, 1, 101].
@@
However, atomic orbital bases generally lack the systematic convergence inherent to other basis types. For instance, the completeness of the plane-wave basis can be controlled by adjusting the kinetic energy cutoff. Even among localized basis sets, there also exist systematically refinable alternatives such as real-space grids [7], truncated spherical waves [31], B-splines [32], Lagrange functions [103], wavelets [24], etc. For atomic orbitals, additional work is needed to establish a well‑defined hierarchy of basis sets with transferable accuracy.
@@
In the past three decades, much interest has been devoted to the development of numerical atomic orbitals (NAOs) [87, 81, 36, 42, 75, 11, 15], whose radial functions are numerically tabulated and are fully flexible within a cutoff radius. In fact, the strategy for generating strictly localized NAOs is not unique. For example, a wide range of methods involve finding the Kohn-Sham orbitals of isolated atoms under suitable confining potentials; these are commonly referred to as pseudo-atomic orbitals, or PAOs. [87, 81, 36, 42, 91, 4]. Based on this idea, Ozaki [75, 74] proposed using PAOs as “primitive functions” to construct NAOs, with coefficients optimized during self-consistent cycles. Alternatively, Blum et al. [11] employed the occupied PAOs as a minimal basis set and expanded it by iteratively selecting the function that most improves a target energy from a predefined pool of candidates. Furthermore, the concept of “correlation consistency” introduced by Dunning [19] has also been investigated in the context of NAOs by Zhang et al. [113]. By minimizing the frozen-core RPA@PBE total energies of free atoms, the resulting basis sets are well suited for correlation methods involving explicit summations over unoccupied states, such as RPA and MP2.
@@
Methods that adopt strategies other than explicit energy optimization also exist. For instance, Sánchez-Portal et al [86, 85] proposed to generate NAOs towards selected reference states by optimizing the “spillage”. In their original scheme, the primitive functions that constitute NAOs are PAOs or Slater-type orbitals, and the reference systems are solids. Following their work, Kenny et al [45] used confined neutral/charged atoms instead of solids as reference systems. Later, Chen, Guo, and He (CGH)  [15] proposed using truncated spherical waves (TSW) as primitive functions to gain greater flexibility and a series of dimers of variable bond lengths as reference systems to improve transferability. Recently, Lin, Ren, and He (LRH) [55] showed that CGH basis sets can be further improved by introducing an extra gradient term to the original spillage function. The above-developed NAOs attain an outstanding balance of efficiency and precision, facilitating accurate large-scale simulations in a wide range of research areas. [49, 23, 80, 92, 70, 14, 64, 79, 59, 57, 13, 51, 71, 47, 69, 73, 76, 93, 97, 39, 52, 9, 21, 54, 84, 61, 60, 62, 94, 40, 114, 115, 66, 96].
Although the NAO basis sets from Ref. 55 deliver remarkable accuracy in calculating structural and electronic properties for numerous molecular and bulk systems, it is still worth studying whether this type of basis set can be further improved. For instance, the hierarchy of these NAO basis sets only extends to standard polarization orbitals, while orbitals with higher angular momentum remain insufficiently explored, which reduces the overall completeness of the basis set. Furthermore, the energy difference between plane waves (PWs) and TSWs with limited angular momentum can reach tens of meV. Since TSWs can be systematically refined toward a complete basis set, this energy gap may be further reduced. [8]. In addition, incorporating more reference states enhances the transferability of NAOs. Meanwhile, plane waves are generally reliable for approaching the complete basis set (CBS). However, spurious interactions with periodic images can trigger artificial bonding across the vacuum. Such issues are difficult to avoid when plane waves are adopted as the expansion basis in practical calculations (Appendix A).
In this work, we propose a scheme to construct NAO basis sets by contracting TSWs, generalizing the spillage minimization scheme via minimizing the kinetic operator trace in the residual space. Our NAOs inherit systematic improvability to approach CBS, with a rigorous strategy for determining key parameters by converging diatomic energy errors against near-complete PWs. Using TSWs as the expansion basis eliminates spurious periodic image interactions, enabling effective transferability enhancement by including virtual states in spillage, which benefits conduction band calculations. Comprehensive benchmarks on molecular and bulk systems verify that the NAOs yield satisfactory precision for key properties like total energy, lattice constant and band gap, and perform well in describing electronic structures.
@@ II.1 Parametrization
The parametrization of the radial shape typically balances several factors, including the computational efficiency of integrals, locality, and flexibility. An ideal form features strict locality and maximum flexibility within a reasonable cutoff radius while allowing for the efficient evaluation of common integrals.
Over the past few decades, several numerical algorithms have been proposed [95, 90, 99, 30] to perform integrations for NAOs efficiently. For strict locality and maximum flexibility, TSWs [31] emerge as a suitable basis for generating NAOs, which gives
$\chi_{l\zeta}(r)=\left\{\begin{matrix}\displaystyle\sum_{q=1}^{N_{l}}j_{l}(\theta_{lq}r/r_{\mathrm{c}})c_{lq\zeta}&r\leq r_{\mathrm{c}}\\[6.0pt] 0&r>r_{\mathrm{c}}\end{matrix}\right.$~~. (1)
Here $j_{l}$ is the spherical Bessel function of the first kind, $\theta_{lq}$ is the $q$-th positive zero of $j_{l}$, $r_{\mathrm{c}}$ is the cutoff radius, and $c_{lq\zeta}$ is the linear combination coefficient. Note that spherical waves are also eigenfunctions of the kinetic energy operator (assuming atomic Rydberg unit)
@@
where $k$ is the norm of momentum. Therefore, similar to the plane wave basis, the number of spherical waves $N_{l}$ can be controlled by a kinetic energy cutoff $E_{\mathrm{c}}$
$N_{l}=\text{max}\{q|\theta_{lq}<r_{c}\sqrt{E_{\mathrm{c}}}\}~~.$ (3)
The full set of TSWs converges to the complete basis set (CBS) as the cutoff radius, maximum angular momentum, and cutoff energy all tend to infinity. For a fixed cutoff radius, constructing an atomic orbital basis is equivalent to determining the contraction coefficients of TSWs.
Note that TSWs do not possess vanishing derivatives at the cutoff radius. To avoid derivative discontinuities, previous works [45, 15] have employed an additional smoothing function. Although this modification has no significant impact on the quality or implementation of NAOs, it not only requires extra effort to determine an optimal smoothing parameter in addition to optimizing the contraction coefficients, but also prevents the orbitals from being pure combinations of TSWs. As a result, the potential advantages offered by certain analytical integral expressions [31, 68] can no longer be exploited.
Rather than modifying the functional form, we propose searching for the subspace where derivatives vanish at the cutoff radius while retaining the original form. Suppose we wish to combine spherical Bessel functions using a set of coefficients $K_{q\lambda}$ to form a new basis set whose first $M$ derivatives vanish at the cutoff $r_{\mathrm{c}}$. We have the following formulas
$\xi_{l\lambda}(r)$ $=\sum_{q=1}^{N_{l}}j_{l}(\theta_{lq}r/r_{\mathrm{c}})K_{q\lambda}~~,$ (4)
@@
Eqs. 4-6 then yield $DK=0$, meaning that the columns of $K$ lie in the null space of $D$. In practice, these can be chosen as the right singular vectors associated with zero singular values. Moreover, since $j_{l}$ satisfies the spherical Bessel equation
@@
it follows that $D_{2q}=(-2/r_{c})D_{1q}$. This implies that any linear combination that suppresses the first derivative automatically eliminates the second derivative as well. If only the first two derivatives are considered, we can construct a set of $N_{l}-1$ basis functions $\{\xi_{l\lambda}\}$ with vanishing first two derivatives from $N_{l}$ spherical Bessel functions. In the remainder of this paper, we refer to $\{\xi_{l\lambda}\}$ as NSW and unless stated otherwise, all orbitals are constructed from NSW with $M=2$.
@@ II.2 Optimization
The contraction of the NSW basis set to obtain optimized NAOs is demonstrated in terms of defining the generalized spillage, choosing reference systems, and the NAO basis set hierarchy. The summarized workflow is shown in Fig. 1.
@@ II.2.1 Generalized Spillage
proposed by Lin et al[55], in which $\hat{p}$ is the momentum operator, we introduce the function $\tilde{\mathcal{S}}$ to be minimized as
@@
This formulation depicts a distinct framework from those of Sánchez-Portal et al., Chen et al. (Eq. 8) and Lin et al. (Eq. 10). Specifically, Eq. 11 corresponds to the trace of a general operator $\hat{O}$ evaluated in the residual subspace spanned by $\{(1-\hat{P}_{\mathbf{k}})|\psi_{n\mathbf{k}}^{\mathrm{NSW}}\rangle\}$. Here, $\{|\psi_{n\mathbf{k}}^{\mathrm{NSW}}\rangle\}$ denotes a set of vectors residing in the space spanned by NSWs, and the subspace spanned by the contracted NSWs (i.e., NAOs) is determined by minimizing as much as possible the residual information defined by the chosen operator.
In this work, $\{|\psi_{n\mathbf{k}}^{\mathrm{NSW}}\rangle\}$ is chosen as the occupied bands, optionally augmented by a subset of unoccupied bands, while $\hat{O}$ is set to the kinetic operator $\hat{p}^{2}$. In contrast, if $|\psi_{n\mathbf{k}}\rangle$ denotes the reference states computed with a plane-wave basis, replacing $\hat{O}$ with the overlap operator $\hat{S}$ or the sum of the overlap and kinetic operators $\hat{S}+\hat{T}$ reduces Eq. 11 to Eq. 8 or Eq. 10, respectively. Since the space spanned by TSWs or NSWs is not strictly a subspace of plane waves due to the real-space truncation of Bessel functions at $r_{\mathrm{c}}$, minimizing the spillage defines a fitting-type problem.
@@ II.2.2 Reference Systems
To ensure the transferability of NAOs, reference systems should cover potential polarization and bonding in real chemical environments. Blum et al. [11] and Chen et al. [15] used homonuclear dimers with various bond lengths. Lin et al. [55] included additional trimers in the generation of triple-zeta level NAOs. Their NAOs all show stable precision across a wide range of systems. In this paper, we use at least 4 bond lengths for each dimer as references. The specific bond lengths are selected so that the corresponding energies relative to equilibrium fall within an energy window of approximately 1.0 eV.
@@ II.3 Basis Set Hierarchy
To systematically approach the maximum angular momentum of NSWs, we use Dunning’s hierarchy [19] to construct NAOs. We define three standard NAO tiers: minimal, polarized valence double-zeta (pVDZ), and polarized valence triple-zeta (pVTZ), together with two variants with constrained angular momentum, labeled pVTZ- and pVQZ=, for comparison.
Our method uses the “shell-wise” optimization scheme developed for the CGH [15] and LRH [55] basis sets, in which basis functions are added and optimized incrementally while keeping all previously generated orbitals fixed. The process starts by constructing a minimal basis corresponding to the valence-electron configuration of the pseudopotential. New functions are then added and optimized with the minimal basis held fixed to form the pVDZ set. In subsequent steps, further functions are added under fixed pVDZ constraints to produce the angular-momentum-restricted triple-zeta basis pVTZ-. Finally, further extensions yield the full triple-zeta pVTZ and the doubly restricted quadruple-zeta pVQZ= basis sets.
@@
Figure 1: Workflow of NAOs generation divided into four steps. First, a plane-wave (PW) basis set is prepared and a set of reference structures is chosen. Second, the superparameters $r_{\mathrm{c}}$ and $l_{\max}$ are chosen such that the energy difference $\epsilon^{\mathrm{PW}}_{\mathrm{NSW}}$ between PW and NSW basis sets is below a threshold. Third, a series of SCF calculations is performed using the converged NSW basis, yield the overlap matrix $S_{\mu\nu}(\mathbf{k})$, the kinetic energy matrix $T_{\mu\nu}(\mathbf{k})$, and the wavefunction coefficients $C_{n\mu}(\mathbf{k})$, where $\mu$ and $\nu$ run over all NSW functions, $n$ denotes eigenstates, and $\mathbf{k}$ denotes different $\mathbf{k}$-points. Finally, the generalized spillage (Eq. 11) is minimized, and the contraction of the NSW basis is performed.